a car travels 10km in 1 hour 12km in next 1hour and 6km in the 3rd 1 hour.calculate the average speed of the car.

Answer by mananth(12270)

total distance traveled 10+12+6=28 km
total time = 3 hours
Average speed = 28/3 = 9.33 km/hr